# A flexible and general semi-supervised approach to multiple hypothesis testing

### v1

@@ Appendix A Supplement
$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1,\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\frac{\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}{P(\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{(1-s)\cdot\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})+1\cdot\mathbb{P}(L_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}$
@@
Lemma S1 essentially states that the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are i.i.d random variables with $\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ independently of all other information: the scores $W$, the side information $\mathbf{x}$, the labels of the false null discoveries, and all the labels $\tilde{L}$.
@@
Recall that the scores $\tilde{W}$ are themselves a function of the values $(\tilde{L},W,x)$ and clearly the one bit p-values are a function of the labels $L_{i}$. Then it follows from the last Lemma that given the scores $\tilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls hypotheses ($i\in N$ with $\tilde{L}_{i}=1$) are i.i.d with $\mathbb{P}(\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})})=\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$, i.e., they stochastically dominate the uniform distribution and hence are valid p-values. Therefore, the condition of Theorem 3 holds so applying, as RESET does, Selective SeqStep+ with $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ to these p-values controls the FDR at level $\alpha$. Note that the original formulation of Selective SeqStep+ is equivalent to Algorithm S1 since we identify the event $\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ with $L_{i}=1$ and $\tilde{p}_{i}=1$ with $L_{i}=-1$. ∎
@@
The proof of Theorem 1 shows that conditional on the learned scores $\tilde{W}$ and the false null labels the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are distributed as i.i.d ±$\pm 1$ RVs, with $\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$. Assuming for a moment that $\mathbb{P}(L_{i}=1)=\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then RESET satisfies the assumption of the multiple decoy version of FDP-SD with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and therefore it controls the FDP at $\alpha$ with confidence $1-\gamma$ (Section 4 of [39]). In the case that $\mathbb{P}(L_{i}=1)<\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then the procedure is conservative with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and the proof in [39] follows without any modification. ∎
@@
We point out some additional minor details regarding our implementation of RESET. We use the randomxForest package, nnet and mgcv to implement the random forest, two-layer neural network and generative additive model in R, respectively. We used the default parameters for the random forest, the default parameters for the two-layer neural network and the default parameters for the generative additive model (except we use drop.intercept = TRUE). If the number of side information variables used by RESET is $\leq 3$, a smoothing spline is fitted with mgcv::s otherwise a natural cubic spline is fitted using 5 degrees of freedom with splines::ns on each side information variable separately (for computational efficiency). For the use of GAM in RESET, we used tryCatch to switch to a random forest in the rare case that the GAM was not able to complete. Some real data sets contain p-values that are ‘zero’. Hence when we calculate $W_{i}=\lvert\Phi^{-1}(p_{i})\rvert$ for such p-values, we obtain Inf in R. We replace instances of Inf for the maximum real-valued score in the dataset. Lastly we used the default settings of RESET for all numerical and real data experiments unless otherwise stated in the main text.
@@
There were some small differences between the implementations in the adaptiveKnockoff package in R and the code used in the original manuscript. Accordingly, we used the updated package version. We also fixed a minor bug, which pruned the candidate set $\mathcal{S}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list. There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g. the initial proportion of hypotheses revealed, reveal_prop. In Simulations 1-3, we used the same parameters set by Ren and Candès from their paper. In Simulation 4, we used the same parameters in Simulation 3.
In the application to peptide detection, the primary scores $W$ may be negative. Since Adaptive Knockoffs encodes the labels $L$ by using the sign of the scores $W$, we shifted our peptide scores so that the minimum score was zero. For AdaKO EM, any peptide score $\leq 10^{-3}$ was subsequently assigned a zero score. Finally we used, reveal_prop = 10% which reveals the hypotheses that are less than or equal to the 10% of non-zero scoring hypotheses (which is the same setting used by the simulations). All other parameters were default.
@@
We implemented AdaPTg using the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define assymetric regions for mirroring the p-values. Accordingly, we followed the same implementation details as the previous section used by AdaPT and set the so-called ‘masking parameters’ as default.
@@
We used the default parameters of zap_asymp from the ZAP package https://github.com/dmhleung/zap. In all simulations and real data experiments, we considered a natural cubic spline basis expansion on each side information variable with six degrees of freedom using splines::ns. This choice was based on their own analysis on two of the real data sets that we consider here, Airway and Bottomly. The exception was in Simulation 6, where we directly used all 100 side information variables rather than applying a spline basis expansion on each of them. In Simulation 5 and 6, we used the test statistics obtained as input, while in the real data experiments we converted the p-values to z-values by using the transformation ±$\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d uniformly.
@@
We implemented AdaFDR using the method.adafdr_test function from the AdaFDR package https://github.com/martinjzhang/adafdr. We note that installation of the package required setting up a separate conda environment so that we could install an older version of Python (v3.6.13). This was because recent versions of Python do not support PyTorch v1.4.0, a requirement of the AdaFDR package setup.py file.
We used the exact same parameters in the vignettes https://github.com/martinjzhang/AdaFDRpaper/tree/master/vignettes to implement AdaFDR on the 10 datasets that Zhang et al. look at. In the gene-drug response data, we used AdaFDR with fast_mode = False. This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when there are few discoveries to be expected or if the number of hypotheses are small (which is not the case for the gene-drug response data).
@@
We downloaded the human proteome (UP000005640, downloaded on 2023/09/23 from UniProt) and prepared a target peptide database along with 5 randomly shuffled decoy peptide databases using Tide-index within the Crux Toolkit v4.1.6809338 [42, 29] with all options set to default. An output containing the target and decoy peptide pairs are conveniently provided using the --peptide-list T option in Tide-index. Each of the RAW 24 HEK293 spectrum files [9] were converted to mzML format using MSConvert 3.0.22314 with the vendor peak-picking filter using the default settings. For each of the 5 decoy databases, we searched each spectrum from the combined 24 mzML spectrum files against the combined target-decoy peptide database using Tide-search [15], using the options --top-match 1 --auto-precursor-window warn --auto-mz-bin-width warn --concat T. The resulting search files were then converted to so-called pin files using the make-pin function in Crux. Each row in the pin files correspond to the optimal peptide-spectrum match (PSM) for each spectrum, the primary score for quantifying this match, called XCorr scores, and a collection of auxilliary information regarding the PSM.
@@
Search algorithms like the one above are typically ‘spectrum-centric’, meaning that for each spectrum, we look for the best matching database peptide, as oppose to the other way round. Consequently, the same database peptide may be matched multiple times with different scores, or have no match at all. Hence for each target peptide, we define $Z_{i}$ as the maximum XCorr score of all the PSMs associated to that target peptide. Similarly for each decoy peptide, we define $\tilde{Z}_{i}$ as the maximum XCorr score of all the PSMs associated to that decoy peptide. For target or decoy peptides that do not appear in the pin file, they get a score of $\infty$- ∞. Then we compete each target and its corresponding paired decoy by recording the winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ along with the label $L_{i}=sign(Z_{i}-\tilde{Z}_{i})$, randomly breaking any ties. Peptides with a winning score of $\infty$- ∞, indicating that neither the target or its paired decoy are in the pin file, are thrown out. We assign each winning peptide the side information defined as the auxiliary information $x_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$. The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table S4.
@@
Running Adaptive Knockoffs on all of the peptide datasets would be infeasible. Therefore, we resorted to estimating its runtimes by considering the time it took to complete a single iteration that determines which hypothesis to reveal next, and multiplying this value by the estimated number of iterations before the procedure stops. Since the time it takes a single iteration depends on the number of hypotheses that were already revealed, we took different such measurements as we varied this number. Specifically, we revealed the labels that were less than or equal to the 10% quantile of non-zero scoring hypotheses, up to a value of $10k$%, in 10% increments. We chose $k$ not to be too large, since hypothetically Adaptive Knockoffs might terminate at the 1% FDR level before it reaches $10k$of the revealed hypotheses. Of course, we are unable to compute when Adaptive Knockoffs might terminate, so we estimated this using the value $r$, the number of target and decoy peptides that scored above the 1% FDR cutoff using RESET ensemble. Then we defined $k=\lfloor(m-r)/m\cdot 10\rfloor$, where $m$ is the total number of hypotheses (peptides) considered by Adaptive Knockoffs. With $\bar{t}_{k}$ being the average of the $k$ measured single-iteration times, we estimated the runtime by $\bar{t}_{k}\cdot(m^{\prime}-r)$, where $m^{\prime}$ is the number of labels that are yet to be revealed (that are above the 10% quantile of non-zero scoring hypotheses) and $m^{\prime}-r$ is our estimate of the number of iterations until Adaptive Knockoffs stop. In Section 5.2.2, we considered 10 repeated applications of RESET ensemble, in which case $r$ is the average of the number of winning target and decoy peptides above the cutoff over each application.
@@
We next consider spectrum files that were searched with variable modifications (---auto-modifications T). Here we need to make a slight adjustment to account for some dependency within the data. Specifically, using variable modifications creates several ‘copies’ of the peptides in the target-decoy database that are distinguished only by slight alterations to the mass of some of the amino acids. As an example, PEPTIDE may generate the following ‘copies’: PE[16]PTIDE, PEPTIDE[16] and PE[16]PTIDE[16], where [16] indicates an increase of 16 Daltons to the amino acid on the left. Consequently, each of these peptides are usually correlated in the data — i.e. if one of these peptides is scored high, than usually they all are. Hence in order to satisfy Assumption 3, we follow the protocol outlined in [18]. That is, we identify all the copies as being the same peptide, ‘PEPTIDE’, with each of the variable modifications ignored. Then we proceed with the same steps as the above paragraph, by determining the maximum score associated with PEPTIDE and recording it as $Z_{i}$ (or $\tilde{Z}_{i}$ if it is a decoy). We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x}_{i})$.
Due to runtime considerations, we used 13 of the smallest narrow search files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript, and only one one of the 10 target-decoy databases. To fairly compare with Adaptive Knockoff, we used the same target-decoy databases for RESET ensemble. All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in the comparison of RESET ensemble and FDP-SD while controlling the FDP.
@@
A standard assumption of the mass spectrometry literature, which is also empirically validated, is that the label $L_{i}$, indicating whether an incorrect target or its decoy scored higher, is uniformly ±$\pm 1$ independently of all other target-decoy pairs and scores [21, 38]. RESET additionally requires this statement to be true if we further condition on the side information. This is obviously true for some side information variables which are constant between the target and corresponding decoys — e.g. PepLen, Charge, lnNumSP, dm, absdM. The other side information variables are score-related — e.g. Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the “standard assumption”. Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied.
S.16 Description and Data preparation using data from Zhang et al.
Each of the 10 datasets from Zhang et al. were downloaded from https://github.com/patrickrchao/AdaPTGMM_Experiments/tree/main/AdaFDR_experiments/data_files, containing the p-value and side information pairs, $(p,\mathbf{x})$. All pairs were then used to report a list of discoveries by each p-value based method with the following exceptions. In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values. These ‘spikes’ concentrate in certain regions of the side information, i.e. when the log-normalized gene counts are low. Hence we removed those p-values with log-normalized gene counts that were less than zero. Figure S1 show the histograms of the p-values before and after this removal.
Figure S1: Histograms of p-values between in the RNA-seq data. The first row of panels plot the histograms before the removal of some of the p-values while the second row of panels plot the histograms after this removal.
@@
Figure S3: Estimated power of each method in numerical simulations. Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction. A description of AdaKO GAM and KO can be found in Section 2.

### v2

@@ 7 Discussion
$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)$ $=\frac{\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)}{P(\tilde{L}_{i}=1)}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{(1-s)\cdot\mathbb{P}(L_{i}=-1)+1\cdot\mathbb{P}(L_{i}=1)}$
$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1)}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}.$
@@
Lemma S1 essentially states that the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are i.i.d. random variables with $\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ independently of all other information: the scores $W$, the side information $\mathbf{x}$, the labels of the false null discoveries, and all the labels $\tilde{L}$.
@@
Recall that the scores $\tilde{W}$ are themselves a function of the values $(\tilde{L},W,x)$ and clearly the one bit p-values are a function of the labels $L_{i}$. Then it follows from the last Lemma that given the scores $\tilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls hypotheses ($i\in N$ with $\tilde{L}_{i}=1$) are i.i.d. with $\mathbb{P}(\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})})=\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$, i.e., they stochastically dominate the uniform distribution and hence are valid p-values. Therefore, the condition of Theorem 3 holds so applying, as RESET does, Selective SeqStep+ with $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ to these p-values controls the FDR at level $\alpha$. Note that the original formulation of Selective SeqStep+ is equivalent to Algorithm S1 since we identify the event $\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ with $L_{i}=1$ and $\tilde{p}_{i}=1$ with $L_{i}=-1$. ∎
@@
The proof of Theorem 1 shows that conditional on the learned scores $\tilde{W}$ and the false null labels the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are distributed as i.i.d. ±$\pm 1$ RVs, with $\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$. Assuming for a moment that $\mathbb{P}(L_{i}=1)=\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then RESET satisfies the assumption of the multiple decoy version of FDP-SD with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and therefore it controls the FDP at $\alpha$ with confidence $1-\gamma$ (Section 4 of [40]). In the case that $\mathbb{P}(L_{i}=1)<\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then the procedure is conservative with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and the proof in [40] follows without any modification. ∎
@@
We point out some additional minor details regarding our implementation of RESET. We use the randomxForest package, nnet and mgcv to implement the random forest, two-layer neural network and generalized additive model in R, respectively. We used the default parameters for the random forest, the default parameters for the two-layer neural network and the default parameters for the generalized additive model (except we use drop.intercept = TRUE). If the number of side information variables used by RESET is $\leq 3$, a smoothing spline is fitted with mgcv::s otherwise a natural cubic spline is fitted using 5 degrees of freedom with splines::ns on each side information variable separately (for computational efficiency). For the use of GAM in RESET, we used tryCatch to switch to a random forest in the rare case that the GAM was not able to complete. Some real datasets contain p-values that are ‘zero’. Hence when we calculate $W_{i}=\lvert\Phi^{-1}(p_{i})\rvert$ for such p-values, we obtain Inf in R. We replace instances of Inf for the maximum real-valued score in the dataset. Lastly we used the default settings of RESET for all numerical and real data experiments unless otherwise stated in the main text.
@@
There were some small differences between the implementations in the adaptiveKnockoff package in R and the code used in the original manuscript. Accordingly, we used the updated package version. We also fixed a minor bug, which pruned the candidate set $\mathcal{S}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list. There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g., the initial proportion of hypotheses revealed, reveal_prop. We used the same parameter settings in Simulation 1 as Simulation 1 in [44], and the same parameter settings in Simulation 2-4 as their Simulation 2 [44].
In the application to peptide detection, the primary scores $W$ may be negative. Since Adaptive Knockoffs encodes the labels $L$ by using the sign of the scores $W$, we shifted our peptide scores so that the minimum score was zero. For AdaKO EM, any peptide score $\leq 10^{-3}$ was subsequently assigned a zero score. Finally we used, reveal_prop = 10% which reveals the hypotheses that are less than or equal to the 10% of positive scoring hypotheses (which is the same setting used by the simulations). All other parameters were default.
@@
We implemented AdaPTg using the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define asymmetric regions for mirroring the p-values. Accordingly, we followed the same implementation details as the previous section used by AdaPT and set the so-called ‘masking parameters’ as default.
@@
We used the default parameters of zap_asymp from the ZAP package https://github.com/dmhleung/zap. In all simulations and real data experiments, we considered a natural cubic spline basis expansion on each side information variable with six degrees of freedom using splines::ns. This choice was based on their own analysis on two of the real datasets that we consider here, Airway and Bottomly. The exception was in Simulation 6, where we directly used all 100 side information variables rather than applying a spline basis expansion on each of them. In Simulation 5 and 6, we used the test statistics obtained as input, while in the real data experiments we converted the p-values to z-values by using the transformation ±$\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d. uniformly.
@@
We applied AdaFDR using the method.adafdr_test function from the AdaFDR package https://github.com/martinjzhang/adafdr. We note that installation of the package required setting up a separate conda environment so that we could install an older version of Python (v3.6.13). This was because recent versions of Python do not support PyTorch v1.4.0, a requirement of the AdaFDR package setup.py file.
We used the exact same parameters in the vignettes https://github.com/martinjzhang/AdaFDRpaper/tree/master/vignettes to implement AdaFDR on the 10 datasets that Zhang et al. look at. Interestingly, the two fMRI datasets only consider the Brodmann label as the side information. We tried using all the available side information, but the statistical power was worse (data not shown). In the gene-drug response data, we used AdaFDR with fast_mode = False. This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when few discoveries are expected or if the number of hypotheses is small (which is not the case for both of the gene-drug response datasets).
@@
We downloaded the human proteome (UP000005640, downloaded on 2023/09/23 from UniProt) and prepared a target peptide database along with 5 randomly shuffled decoy peptide databases using Tide-index within the Crux Toolkit v4.1.6809338 [43, 29] with all options set to default. An output containing the target and decoy peptide pairs are conveniently provided using the --peptide-list T option in Tide-index. Each of the RAW 24 HEK293 spectrum files [9] were converted to mzML format using MSConvert 3.0.22314 with the vendor peak-picking filter using the default settings. For each of the 5 decoy databases, we searched each spectrum from the combined 24 mzML spectrum files against the combined target-decoy peptide database using Tide-search [15], using the options --top-match 1 --auto-precursor-window warn --auto-mz-bin-width warn --concat T. The resulting search files were then converted to so-called pin files using the make-pin function in Crux. Each row in the pin files correspond to the optimal peptide-spectrum match (PSM) for each spectrum, the primary score for quantifying this match, called XCorr scores, and a collection of auxiliary information regarding the PSM.
@@
Search algorithms like the one above are typically ‘spectrum-centric’, meaning that for each spectrum, we look for the best matching database peptide, as oppose to the other way round. Consequently, the same database peptide may be matched multiple times with different scores, or have no match at all. Hence for each target peptide, we define $Z_{i}$ as the maximum XCorr score of all the PSMs associated to that target peptide. Similarly for each decoy peptide, we define $\tilde{Z}_{i}$ as the maximum XCorr score of all the PSMs associated to that decoy peptide. For target or decoy peptides that do not appear in the pin file, they get a score of $\infty$- ∞. Then we compete each target and its corresponding paired decoy by recording the winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ along with the label $L_{i}=sign(Z_{i}-\tilde{Z}_{i})$, randomly breaking any ties. Peptides with a winning score of $\infty$- ∞, indicating that neither the target or its paired decoy are in the pin file, are thrown out. We assign each winning peptide the side information defined as the auxiliary information $\mathbf{x}_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$. The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table S4.
@@
Running Adaptive Knockoffs (AdaKO) on all peptide datasets would be infeasible. Therefore, we resorted to estimating its runtimes by considering the time it took to complete a single iteration that determines which hypothesis to reveal next, and multiplying this value by the estimated number of iterations before the procedure stops.
Since the time it takes a single iteration depends on the number of hypotheses that were already revealed, we took different such measurements as we varied this number, and averaged them. Specifically, as AdaKO does, we first revealed the labels that were scored less than or equal to the 10% quantile of non-zero scoring hypotheses, and recorded the time it takes for a single interation from this point. To be clear, the revealed labels thus includes those zero-scoring hypotheses since they will be less than this score cutoff — it is only the 10% quantile that considers the non-zero scoring hypotheses. Then we took similar measurements by increasing the ‘10%’ value, in 10% increments, up to a value of ‘10k%’ where we define $k$ next.
We chose the value of $k$ so that the proportion of revealed labels, $10k\%$, estimates the number of revealed labels when AdaKO terminates at the 1% FDR level. That way, the $k$ times should provide an accurate sample of the number single iteration times across all of AdaKO’s iterations until it terminates. Of course, we are unable to compute when Adaptive Knockoffs might terminate, so we estimated this using the value $q$, the number of target and decoy peptides that scored above the 1% FDR cutoff using RESET Ensemble. Then we defined $k=\lfloor(m-q)/m\cdot 10\rfloor$, where $m$ is the total number of hypotheses (peptides) considered by Adaptive Knockoffs. We estimated the runtime by
$\bar{t}_{k}\cdot(m^{\prime}-q),$
where $\bar{t}_{k}$ is the average of the $k$ measured single-iteration times, and $m^{\prime}$ is the number of labels that are yet to be revealed at the start of AdaKO’s iterations, i.e., the number of labels with scores above the 10% quantile of non-zero scoring hypotheses. Thus the value $m^{\prime}-q$ is our estimate of the number of iterations until Adaptive Knockoffs stop.
In Section 5.2.2, we considered 10 repeated applications of RESET Ensemble, varying RESET Ensemble’s internal seed. In this case, we define $\hat{\ell}$ as the average of the number of winning target and decoy peptides above the cutoff over each application.
@@
We next consider spectrum files that were searched with variable modifications (---auto-modifications T). Here we need to make a slight adjustment to account for some dependency within the data. Specifically, using variable modifications creates several ‘copies’ of the peptides in the target-decoy database that are distinguished only by slight alterations to the mass of some of the amino acids. As an example, PEPTIDE may generate the following ‘copies’: PE[16]PTIDE, PEPTIDE[16] and PE[16]PTIDE[16], where [16] indicates an increase of 16 Daltons to the amino acid on the left. Consequently, each of these peptides are usually correlated in the data — i.e., if one of these peptides is scored high, than usually they all are. Hence in order to satisfy Assumption 3, we follow the protocol outlined in [18]. That is, we identify all the copies as being the same peptide, ‘PEPTIDE’, with each of the variable modifications ignored. Then we proceed with the same steps as the above paragraph, by determining the maximum score associated with PEPTIDE and recording it as $Z_{i}$ (or $\tilde{Z}_{i}$ if it is a decoy). We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x}_{i})$.
Due to runtime considerations, we used 13 of the smallest narrow search files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript, and only one of the 10 target-decoy databases. To fairly compare with Adaptive Knockoffs, we used the same target-decoy databases for RESET ensemble. All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in the comparison of RESET ensemble and FDP-SD while controlling the FDP.
@@
A standard assumption of the mass spectrometry literature, which is also empirically validated, is that the label $L_{i}$, indicating whether an incorrect target or its decoy scored higher, is uniformly ±$\pm 1$ independently of all other target-decoy pairs and scores [21, 39]. RESET additionally requires this statement to be true if we further condition on the side information. This is obviously true for some side information variables which are constant between the target and corresponding decoys — e.g., PepLen, Charge, lnNumSP, dm, absdM. The other side information variables are score-related — e.g., Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the ‘standard assumption’. Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied.
S.16 Preparation of data with p-values
Each of the 10 datasets from Zhang et al. were downloaded from https://github.com/patrickrchao/AdaPTGMM_Experiments/tree/main/AdaFDR_experiments/data_files, containing the p-value and side information pairs, $(p,\mathbf{x})$. All pairs were then used to report a list of discoveries by each p-value based method with the following exceptions. In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values. These ‘spikes’ concentrate in certain regions of the side information, i.e., when the log-normalized gene counts are low. Hence we removed those p-values with log-normalized gene counts that were less than zero. Figure S1 show the histograms of the p-values before and after this removal. The gene-drug response data was obtained from the adaptMT package.
Figure S1: Histograms of p-values in the RNA-seq data before and after p-value removal. Top row: the original histograms of the p-values. Bottom row: the corresponding histograms after removal of the p-values associated with low counts.
@@
Figure S3: Estimated power of each method in numerical simulations. Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction. A description of the AdaKO methods and KO can be found in Section 2.